If n=15 and p=.4, then the standard deviation of the binomial distribution is
Answer #1

Standard deviation of binomial distribution is Sqrt(np(1-p))

n = 15, p = 0.4

np(1-p) = 15*0.4*0.6 = 3.6

Standard deviation of binomial distribution is 1.8974